# PyTorchGeoNodes: Enabling Differentiable Shape Programs for 3D Shape Reconstruction

### v1

@@ Abstract
We propose PyTorchGeoNodes, a differentiable module for reconstructing 3D objects from images using interpretable shape programs. In comparison to traditional CAD model retrieval methods, the use of shape programs for 3D reconstruction allows for reasoning about the semantic properties of reconstructed objects, editing, low memory footprint, etc. However, the utilization of shape programs for 3D scene understanding has been largely neglected in past works. As our main contribution, we enable gradient-based optimization by introducing a module that translates shape programs designed in Blender, for example, into efficient PyTorch code. We also provide a method that relies on PyTorchGeoNodes and is inspired by Monte Carlo Tree Search (MCTS) to jointly optimize discrete and continuous parameters of shape programs and reconstruct 3D objects for input scenes. In our experiments, we apply our algorithm to reconstruct 3D objects in the ScanNet dataset and evaluate our results against CAD model retrieval-based reconstructions. Our experiments indicate that our reconstructions match well the input scenes while enabling semantic reasoning about reconstructed objects.
              Shape parameter                                 Differentiable graph         Generated 3D shape
Our reconstructions
Figure 1: Our PyTorchGeoNodes is a framework that enables differentiable shape programs in PyTorch. Given interpretable input shape parameters, a shape program generates a mesh for a 3D object. Our differentiable pipeline makes possible the flow of gradients from the generated shape to the continuous parameters. We show a variety of sofas and armchairs generated by a single shape program. The method we propose can retrieve these shapes using a sequence of RGB-D scans as input, estimating both the continuous parameters (Width, Depth, etc.) and the discrete parameters (Has Left Arm, Is L-Shaped, etc.).
@@ 1 Introduction
In 3D scene understanding, many different representations of 3D objects that can be recovered from images have already been considered: voxels [29, 20], meshes [12, 38], SDFs [24], NeRFs [7], Gaussian Splatting [34], exemplars from a shape database [3]. SDFs, NeRFs, or Gaussian Splatting can generalize to shapes unseen during training but can still produce artifacts for example when parts of the object are occluded in scenes captured “in the wild”. Moreover, their memory footprint can become large. Using exemplars guarantees that the retrieved shapes look good as they are usually created by human designers, however the database may not contain shapes similar to the target object.
In this paper, we investigate the use of interpretable shape programs to represent objects. Shape programs are a form of procedural modeling [11] introduced in [35] for representing 3D objects. The top of Figure 1 illustrates a shape program generating sofa and armchair shapes.
@@
Shape programs are a computational graph of parameterized geometric operations that generate various 3D shapes from an object category given input shape parameters, such as object dimensions as well as “semantic” properties like ’with armrests’. Thanks to this procedural modeling approach, they can capture all the shape variation within an object category so that the generated shapes remain appealing and can fit many different instances. By design, they also allow the decomposition of generated shapes into smaller shape primitives. This makes possible reasoning about semantic components, which is an advantage over more traditional CAD model retrieval methods. Other nice properties include being easily editable and having a small memory footprint. They became a popular tool for building reconstruction [39, 23, 18, 36] and the interest around shape programs for objects is growing [35, 15, 16, 26]. Blender [9] already provides support for them. A research challenge targeting reconstruction of walls, doors, and windows using such structured languages was organized by Meta in 2024 111[https://eval.ai/web/challenges/challenge-page/2115/overview (Accessed: March 2024)].
@@
There have been a few attempts in the computer vision community to use procedural shapes [30, 26]. Recovering the parameters of a shape program from an image is however very challenging. Typically, a shape generated by a shape program depends on a number of parameters. Some parameters such as the width and height of the objects take continuous values but other parameters take only discrete values, for example the number of shelves in a bookshelf. Optimizing on these parameters turned out to be tricky: For example, changing the number of shelves from 5 to 6 changes significantly the shape and appearance of a bookshelf.
@@
We introduce two contributions to make possible the estimation of shape programs’ parameters from images with a render-and-compare algorithm. The first contribution is a “compiler” that generates PyTorch [25] code for a shape program, which can be defined in Blender for example. This code generates 3D shapes that are differentiable with respect to the continuous parameters and can therefore be used easily with 3D reconstruction algorithms to optimize the continuous parameters.
While our compiler can be used to optimize over the continuous parameters, there remains the challenge of estimating the discrete parameters. Our experiments show that random sampling is highly inefficient. In a simple experiment of identifying parameters of a sofa, random search needed $41$ minutes to find an acceptable solution. Our second contribution is to show how to adapt recent methods based on Monte Carlo Tree Search to our problem of jointly optimizing the continuous and the discrete parameters of a shape program. For the same example, our method finds a better solution in $2.3$ minutes.
We evaluate our method on both synthetic and real scenes. In addition to the usual shape metrics, we also evaluate how well we retrieve the shape program discrete parameters. Because existing datasets are not annotated with the value of these parameters, we manually annotated 276 objects. In summary, our main contribution is a general tool to create PyTorch code from Blender shape programs. In addition, we provide a method for estimating the parameters of shape programs from views together with the poses of the objects. We will make our code available on our project page, https://vevenom.github.io/pytorchgeonodes/, with the hope that our compiler will encourage the use of shape programs and that our method will inspire other researchers to work on recovering shape programs from images. Beyond shape retrieval, differentiable rendering of shape programs could open the door to methods creating shape programs from images in a self-supervised way, an exciting research direction.
@@ 2.2 Shape Programs
Procedural modelling is an attractive way of generating 3D shapes in computer graphics, and as such it is an interesting feature of different 3D modelling software. In general, procedural modelling is a family of techniques that generate 3D content by abiding to a set of rules. In machine learning, [35] introduced a procedural language that enables the generation of a variety of shapes in form of so called shape programs. In [35], shape programs are written in Domain Specific Language (DSL) for shapes with two types of statements. “Draw” statements describes shape primitive, such as cuboid or a cylinder, together with its geometric and semantic properties. “For” statements represent loops of sub-programs including parameters that indicate how the sub-program should be repeatedly executed. [35] also showed that it is possible to learn to infer shape programs that describe input 3D shapes. However, as it is based on basic geometric primitives, it becomes difficult to model more complex geometries, and resulting shapes can be non-representative of the target category. Other methods suffer from the same limitations [15, 16].
@@
GeoCode [26] recently introduced hand-designed shape programs as a mapping between 3D shapes of a specific object category and human-interpretable parameter space. The language used for developing shape programs is based on the geometry node feature of Blender, an open-source modelling software, which, in contrast to [35], allows designing of more expressive shape programs that can represent and interpolate between a variety of shapes by simply modifying human-interpretable program’s parameters. Some nodes define a shape primitive or pre-designed more complex shapes, e.g., chair seat or chair leg. Other nodes represent geometric transformations on individual shapes, or how two shapes should be connected, e.g., legs should always be connected to the bottom of the seat. There are nodes representing mathematical operations such as addition and multiplication, and there are nodes that are used to replicate specific shapes, e.g., in case of chair legs we can add a replicator node to create four instances of a chair leg of the same shape. There are also nodes that serve as control statements, e.g., it is possible to switch between a four-legged chair and a one-legged chair. A pre-defined set of input program’s parameters allow non-expert user to easily control shape-generation process. [26] also showed that it is possible to learn the mapping from input point clouds of 3D shapes, or a sketch, to program’s parameters.
3 PyTorchGeoNodes: Differentiable Computational Graphs for Shape Programs
@@
In this section, we discuss shape programs and how they can be beneficial for tasks in 3D scene understanding. We then describe our framework for enabling efficient gradient-based optimization for shape programs and, in addition, allowing to translate shape programs designed in Blender to PyTorch code to enable differentiable optimization.
In Section 4, we explain how our framework can be applied for reconstructing 3D objects from images.
3.1 Shape Programs as Computational Graphs
@@
  • An ’Input’ node contains input parameters of the shape programs and their value ranges. We consider floating, integer, and boolean values parameters;
  • A ’Math’ node performs mathematical operations on given input such as addition, subtraction, multiplication, division, and comparison;
@@
  • A ’Primitive’ node creates a geometric primitive, e.g. a cuboid or a cylinder;
  • A ’Transform’ node applies scaling, rotation, and translation from input vectors to the input geometry;
@@
  • A ’Join geometry’ node merges different geometries;
@@
Note that for example, the Transform node depends on continuous parameters, the Points on instances node on integer parameters, and the Switch node on boolean parameters.
Figure 2: Gradient-based optimization of continuous parameters of a shape program for the sofa category. From an initial estimate of the parameters of the object, we can perform gradient descent on the parameters based on a 3D geometric loss term. In contrast to methods that directly optimize the reconstructed mesh, PyTorchGeoNodes allows optimization in the parameter space which has several benefits. From the resulting shapes in this example, it is observable that individual parameters can be scaled independently targeting only specific parts of the shape geometry while preserving the compactness of the 3D shape at the same time.
3.2 Blender Geometry Nodes to PyTorchGeoNodes
Figure 3: A shape program in Blender generating dividing boards of a cabinet. The computational graph is designed and visualized in Blender using Geometry Nodes feature. Underneath, we show how we abstract the different nodes using PyTorch tensors and PyTorch3D meshes. The input node takes input parameters, here {Width: 0.5, Dividing Board Thickness: 0.04, Height: 0.6, Number of Dividing Boards: 5, Board Thickness: 0.04} and feeds them to a series of operations. The blue nodes are arithmetic and concatenation nodes, which transform input parameters and feed the results to geometry nodes, in green. In this example, we generate a cuboid mesh and instantiate a line of points which generates the final geometry for dividing boards. In practice, this shape program is part of a larger shape program for modeling cabinets.
Our framework provides different computational nodes that reimplement the functionalities of the individual geometry nodes in Blender. We visualize this in Figure 3. More exactly, for every node type in Blender, we implement a corresponding node type with the same functionalities using PyTorch, or PyTorch3D[27] in case of geometric operations. We note that shape programs from Blender are already saved as a graph structure which simplifies the translation process. Therefore, our framework is designed to enable automatic translation from shape programs designed in Blender into PyTorch code and we can design shape programs directly using the Blender interface.
To make our implementation computationally efficient, we considered different aspects. First, the nodes are ordered using a topological sort so that during the forward pass of a node, all inputs are readily available, which avoids the need for recursive calls. In addition, we cache outputs of nodes that do not depend on input shape parameters such that they do not need to be recomputed in subsequent forward calls. We provide additional implementation details in the supplementary material. While computation time depends on the complexity of shape programs, in our experiments inference time is always less than $20$ms.
Figure 2 shows a simple example. Given a shape program created in Blender, our parser generates a differentiable PyTorchGeoNodes code. We can then use this code in PyTorch in an optimization procedure where we fit shape parameters to generate a shape consistent with the input scene.
@@ 4 Search Algorithm
Figure 4: Reconstructing 3D objects with our PyTorchGeoNodes and our method. At every iteration, our MCTS-based search algorithm first selects a set of shape parameters and object pose. These are then passed through the differentiable computational graph, parsed by our PyTorchGeoNodes framework for a shape program, that generates a 3D mesh based on the selected shape parameters. Then, we compute the loss that evaluates how well the shape fits the input scene. We further optimize the continuous parameters of our shape program using gradient descent and update the search tree based on this loss.
We provide in this section a way to retrieve the parameters of a shape program for a given object from several RGB-D scans of this object. We do this by exploiting our PyTorchGeoNodes framework. Following many current works, we use a render-and-compare approach as it appears to provide the best results [2]. More exactly, we look for shape parameters ${\cal P}^{*}$ by minimizing the following objective function:
${\cal L}({\cal P})=$ $\sum_{i}\big{(}|D_{i}-M_{D}(D^{r}_{i}({\cal P}),D_{i},S_{i})|+|N(D_{i})-N(M_{D}(D^{r}_{i}({\cal P}),D_{i},S_{i}))|\big{)}\;+$
$\lambda_{\text{CD}}{\text{CD}}({\text{PC}},{\text{Sh}}({\cal P}))\>.$ (1)
$D_{i}$ and $N_{i}$ denote the depth map and normal map respectively of the $i$-th view, and $S_{i}$ is the estimated segmentation of object per input view, obtained by running a segmentation network on the input color image of the $i$-th view, as explained in Section 5.2. $D^{r}_{i}({\cal P})$ is the depth map of the shape generated by the shape program given parameters ${\cal P}$ corresponding to the $i$-th view. To deal with occlusions of the target object in the input views, we merge the observed depth map $D_{i}$ with the predicted depth map $D^{r}_{i}({\cal P})$ with operator $M_{D}$: For each pixel where $S_{i}=1$, $M_{D}$ returns the values of $D^{r}_{i}({\cal P})$. For each of the other pixels, $M_{D}$ returns the minimum depth between $D_{i}$ and $D^{r}_{i}({\cal P})$ at this pixel. $N(D)$ returns the normal map for depth map $D$.
We notice that including the Chamfer distance ${\text{CD}}({\text{PC}},{\text{Sh}}({\cal P}))$, between the scene and the generated shape improves the speed of convergence. This was observed in [1] as well. This can be explained by the fact that the Chamfer distance has a larger radius of convergence than the render-and-compare terms and can be used to guide the optimization. PC represents the point cloud reconstructed from the depth maps, ${\text{Sh}}({\cal P})$ is the shape generated by the shape program, and CD denotes the Chamfer distance of the point cloud to the generated shape. It is weighted by $\lambda_{\text{CD}}=10$ based on our empirical evaluations. Note that this objective function is not part of our contribution and our framework could handle any other function that is differentiable.
We consider object pose as additional continuous parameters in ${\cal P}$ to enable the rendering of shape programs into scene views. In practice, we assume an initialization of object translation is given and we are looking for rotation around the y-axis only as we explain in Section 5.
We use our PyTorchGeoNodes framework to compute the $D^{r}_{i}({\cal P})$, $N^{r}_{i}({\cal P})$, and ${\text{Sh}}({\cal P})$. Thanks to this framework, PyTorch can thus compute their derivatives and the derivatives of ${\cal L}$ with respect to the continuous parameters in ${\cal P}$.
However, some parameters in ${\cal P}$ are discrete, i.e., integers or booleans. Therefore, we cannot optimize ${\cal L}$ using a simple gradient descent. Also, ${\cal L}$ has no special form that we can exploit to optimize it. We therefore resort to an approach related to the one in [2], which combines stochastic search, more specifically Monte Carlo Tree Search (MCTS), and gradient descent which is enabled by our PyTorchGeoNodes framework.
4.1 MCTS for Shape Parameter Search
Monte Carlo Tree Search (MCTS) is a stochastic search algorithm that relies on guided tree search and Monte Carlo simulations to maximize a given scoring function. A major advantage of MCTS is that it can explore the most promising parts of the tree simply based on the scoring function. It can also avoid local minima and can shift the search focus on less explored sub-trees. It was used in [33] to solve several scene understanding problems, and more details can be found in [33] or in [6] which is a more general survey. We show here how to adapt it to our problem. Further details about different hyper-parameters for our method are given in the supplementary material.
Shape parameter tree. We organize the shape parameters in ${\cal P}$ into a tree structure, with nodes at each level of the tree corresponding to a range of values for a specific shape parameter. In the case of continuous values, we first discretize them by linearly sampling values within a valid range.
In principle, we could only use gradient-based optimization to find the optimal values for the continuous parameters. However, even for these parameters, the objective function ${\cal L}$ is difficult to optimize when no good initial guess is available and gradient descent can lead to poor local minima. Instead, our search algorithm combines continuous and discrete optimization to determine values of continuous parameters. Therefore, every path in the tree instantiates parameters that can be fed to our shape program to generate a 3D shape. We define the order of shape parameters tree based on how much individual parameters influence the shape of generated objects which is further discussed in supplementary material. Therefore, this guarantees that objects in deeper parts of the tree become geometrically more consistent with each other.
Modified selection criterion. At every iteration, MCTS typically selects nodes ${\cal N}$ that maximize the Upper Confidence Bound (UCB):
${\text{UCB}}({\cal N})={\text{score}}({\cal N})+\lambda\sqrt{\frac{\ln{\#{\text{visits}}}({\text{parent}}({\cal N}))}{\#{\text{visits}}({\cal N})}}\>,$ (2)
where ${\text{score}}({\cal N})$ is a score for node ${\cal N}$ which is accumulated over evaluations of a negative of our objective function ${\cal L}({\cal P})$ for the past iterations when node ${\cal N}$ was selected. #$\#{\text{visits}}({\cal N})$ is the number of visits to node ${\cal N}$ in previous iterations. We set $\lambda=0.2$ in all our experiments to weight the exploration term.
In our case, each node ${\cal N}$ is associated to a pair of a parameter $p$ in ${\cal P}$ and a discretized value $v$ for this parameter. Within the tree, many nodes will be associated to a given pair $(p,v)$. In the original formulation of MCTS, these nodes remain independent. For our problem however, it is interesting to make the nodes that correspond to the same pair $(p,v)$ share the same score. This is because assigning value $v$ to $p$ will have a similar influence on the objective function ${\cal L}({\cal P})$ regardless of the values assigned to the other parameters in ${\cal P}$.
We thus store the scores not for each node but for each pair $(p,v)$ and modify the computation of UCB into:
${\text{UCB}}({\cal N})={\text{score}}({\text{pv}}({\cal N}))+\lambda\sqrt{\frac{\ln{\#{\text{iterations}}}}{\#{\text{visits}}({\text{pv}}({\cal N}))}}\>,$ (3)

### v2

@@ Abstract
We propose PyTorchGeoNodes, a differentiable module for reconstructing 3D objects and their parameters from images using interpretable shape programs. Unlike traditional CAD model retrieval, shape programs allow reasoning about semantic parameters, editing, and a low memory footprint. Despite their potential, shape programs for 3D scene understanding have been largely overlooked. Our key contribution is enabling gradient-based optimization by parsing shape programs, or more precisely procedural models designed in Blender, into efficient PyTorch code. While there are many possible applications of our PyTochGeoNodes, we show that a combination of PyTorchGeoNodes with genetic algorithm is a method of choice to optimize both discrete and continuous shape program parameters for 3D reconstruction and understanding of 3D object parameters. Our modular framework can be further integrated with other reconstruction algorithms, and we demonstrate one such integration to enable procedural Gaussian splatting. Our experiments on the ScanNet dataset show that our method achieves accurate reconstructions while enabling, until now, unseen level of 3D scene understanding.
Jointly recovered geometry and parameters of target objects from RGB-D scans
Refer to caption Refer to caption Refer to caption
Views of Target object Recovered Gaussians Procedural Gaussians editing
Figure 1: Our PyTorchGeoNodes is a framework that enables differentiable shape programs for 3D reconstruction. Our semantically-rich reconstructions of 3D objects from RGB-D scans enable understanding of important metric parameters of individual parts for different objects. These parameters are physically bound to the 3D reconstruction as PyTorchGeoNodes enables flow of gradients from the generated shape to the parameters of the object. Our PyTorchGeoNodes is modular, e.g. we can use it to enable procedural modeling for Gaussian splatting. In the example, our method recovers Gaussians for the target sofa and ’clones’ all details from the visible armrest to the occluded armrest in the scene. Our integration enables procedural editing of Gaussians and we can alter shape parameters to edit the reconstruction: e.g. by extending or shrinking width of the seat, removing backrest, left armrest, changing height of the backrest and depth of the seat etc.
@@ 1 Introduction
Various 3D object representations have already been considered for 3D scene understanding: voxels [31, 21], meshes [11, 38], SDFs [25], NeRFs [5], Gaussian Splatting [34], exemplars from a shape database [3]. SDFs, NeRFs, or Gaussian Splatting generalize to shapes unseen during training but produce artifacts for example when parts of the object are occluded in scenes captured “in the wild”. Their memory footprint can also become large. Using exemplars guarantees that the retrieved shapes look good as they are created by human designers, however the database may not contain shapes similar to the target object.
In this paper, we investigate the use of interpretable shape programs to represent objects. Shape programs are a form of procedural modeling [10] introduced in [35] for representing 3D objects. While shape programs have already been considered for 3D scene understanding [32, 27], their use in this context remains largely unexplored. As shown in Figure 1, our method can be used to accurately retrieve geometry and 3D parameters of target objects, together with its texture even in presence of partial occlusions.
Shape programs can be seen as computational graphs of parametrized geometric operations that generate various 3D shapes from an object category given input shape parameters, such as object dimensions as well as “semantic” properties like existence of armrest or backrest in case of chairs. Multiple similar parts, for example the legs of a chair or the shelves of a bookshelves—are obtained by first generating a single instance, duplicating it and applying geometric transformations such as rotations and translations as needed.
@@
Thanks to this procedural modeling approach, shape programs generate shapes that remain appealing and that can fit many different instances. Alongside being interpretable, the shapes they generate are editable in an intuitive manner and have a very small memory footprint. In fact, they are already a popular tool for building reconstruction [39, 24, 19, 37] and the interest around shape programs for objects is growing [35, 15, 16, 27]. Blender [8] already provides support for them. A research challenge targeting reconstruction of walls, doors, and windows using such structured languages was organized by Meta in 2024 111[https://eval.ai/web/challenges/challenge-page/2115/overview (Accessed: November 2024)].
@@
Recovering parameters of a shape program from an image is however challenging. Typically, a shape generated by a shape program depends on a number of parameters. Some parameters such as the width and height of the objects take continuous values but other parameters take only discrete values, e.g. the number of shelves in a bookshelf.
Optimizing on these parameters turns out to be tricky: For example, changing the number of shelves from 5 to 6 changes significantly the shape and appearance of a bookshelf. There have been a few attempts in the computer vision community to use procedural shapes [32, 27]. In [27] the authors train a neural network to predict the parameters of a shape program. While this method shows impressive results using synthetic data, we show in Section 5.3 that on real data such an approach fails to generalize well, due to the lack of real world image data annotated with shape parameters. Using synthetically generated data to train such models also performs very badly: While shape programs capture very well the global structure of objects, they do not represent details. As a result, the model can hardly generalize from synthetic images to real ones. In comparison, our proposed method utilizes an objective function which is not based on learning and does not need training data, hence it is able to generalize well to real world scenarios.
We introduce three main contributions which enable the estimation of shape programs’ parameters from images:
@@
  • Our first contribution is a ‘compiler‘ that generates PyTorch [26] code for a shape program, which can be defined in Blender for example. This code generates 3D shapes that are differentiable with respect to the continuous parameters and can therefore be used easily with 3D reconstruction algorithms to optimize the continuous parameters.
  • Second, we contribute a method to obtain efficiently and reliably all shape program parameters from RGB-D scans: While our compiler can be used to optimize over the continuous parameters, the challenge of estimating the discrete parameters remains. We propose to adapt methods based on genetic algorithms to our problem of jointly optimizing the continuous and the discrete parameters of a shape program.
  • Our third contribution is an extension of shape programs to capture fine details and appearance of objects. We do this by augmenting our programs with Gaussians. Once we retrieved the shape parameters, we initialize our procedural Gaussians over the primitives and optimize them as in Gaussian splatting [17]. The key difference with regular Gaussian Splatting is that our Gaussians are constrained by the procedural model: The Gaussians sampled on ’cloned’ base primitives are tied together. e.g. as the legs of a chair are ’clones’ of the same base primitives, the occluded part of a leg can still be recovered if it is visible for another leg.
We evaluate our method on both synthetic and real scenes. In addition to the usual metrics, we also evaluate how well we reconstruct the shape program continuous and discrete parameters. Because existing datasets are not annotated with the value of these parameters, we manually annotated $176$ objects.
In summary, our main contribution is a general tool to create PyTorch code from Blender shape programs. In addition, we provide a method for estimating the parameters of shape programs from views together with the poses of the objects. Our PyTorchGeoNodes is a modular tool and it can be extended to novel use cases in 3D scene understanding. The source code is available to the general public.
@@ 2.2 Shape Programs
Procedural modelling is an attractive way of generating 3D shapes in computer graphics, and as such it is an interesting feature of different 3D modelling software. In general, procedural modelling is a family of techniques that generate 3D content by abiding to a set of rules. A variety of methods were proposed that combine procedural modeling with machine learning approaches, summarized in detail in [30]. For example, [35] introduced a procedural language that enables the generation of a variety of shapes in form of so called shape programs. In [35], shape programs are written in Domain Specific Language (DSL) for shapes with two types of statements. “Draw” statements describes shape primitive, such as cuboid or a cylinder, together with its geometric and semantic properties. “For” statements represent loops of sub-programs including parameters that indicate how the sub-program should be repeatedly executed.
[35] also showed that it is possible to learn to infer shape programs that describe input 3D shapes. However, as it is based on basic geometric primitives, it becomes difficult to model more complex geometries, and resulting shapes can be non-representative of the target category. Other methods suffer from the same limitations [15, 16].
@@
GeoCode [27] recently introduced hand-designed shape programs as a mapping between 3D shapes of a specific object category and human-interpretable parameter space. The language used for developing shape programs is based on the geometry node feature of Blender, an open-source modeling software, which, in contrast to [35], allows designing of more expressive shape programs that can represent and interpolate between a variety of shapes by simply modifying human-interpretable program’s parameters.
Figure 2: A procedural model in Blender generating dividing boards of a cabinet. The computational graph is designed and visualized in Blender using Geometry Nodes feature. Underneath, we show how we abstract the different nodes using PyTorch tensors and PyTorch3D meshes. The input node takes input parameters, here {Width: 0.5, Dividing Board Thickness: 0.04, Height: 0.6, Number of Dividing Boards: 5, Board Thickness: 0.04} and feeds them to a series of operations. The blue nodes are arithmetic and concatenation nodes, which transform input parameters and feed the results to geometry nodes, in green. In this example, we generate a cuboid mesh and instantiate a line of points which generates the final geometry for dividing boards. In practice, this shape program is part of a larger shape program for modeling cabinets.
Some nodes define a shape primitive or pre-designed more complex shapes, e.g., chair seat or chair leg. Other nodes represent geometric transformations on individual shapes, or how two shapes should be connected, e.g., legs should always be connected to the bottom of the seat. There are nodes representing mathematical operations such as addition and multiplication, and there are nodes that are used to replicate specific shapes, e.g., in case of chair legs we can add a replicator node to create four instances of a chair leg of the same shape. There are also nodes that serve as control statements, e.g., it is possible to switch between a four-legged chair and a one-legged chair. A pre-defined set of input program’s parameters allow non-expert user to easily control shape-generation process. [27] also showed that it is possible to learn the mapping from input point clouds of 3D shapes, or a sketch, to program’s parameters.
3 PyTorchGeoNodes
@@
In this section, we discuss shape programs and how they can be beneficial for tasks in 3D scene understanding. We then describe our framework for enabling efficient gradient-based optimization for shape programs and, in addition, allowing to translate shape programs designed in Blender to PyTorch code to enable differentiable optimization. In Section 4, we explain how our framework can be applied for reconstructing 3D objects from images.
@@
  • An ’Input’ node contains input parameters of the shape programs and their value ranges. We consider floating, integer, and boolean values;
  • A ’Math’ node performs mathematical operations on its input (addition, subtraction, multiplication, division, and comparison);
@@
  • A ’Primitive’ node creates a geometric primitive, e.g., a cuboid or a cylinder;
  • A ’Transform’ node applies scaling, rotation, and translation to the input geometry;
@@
  • A ’Join geometry’ node merges geometries;
@@
Note that for example, the Transform node depends on continuous, the Points on instances node on integer, and the Switch node on boolean.
Our framework provides differentiable computational nodes that reimplement the functionalities of the individual geometry nodes in Blender as we show in Figure 2. More exactly, for every node type in Blender, we implement a corresponding node type with the same functionalities using PyTorch, or PyTorch3D [29] in case of geometric operations. In practice, we designed shape programs directly using the Blender interface. Given a shape program created in Blender, our parser generates a differentiable PyTorchGeoNodes code. We can then use this code in PyTorch in an optimization procedure where we fit shape parameters to generate a shape consistent with the input scene.
In the supplementary material, we detail our design choices and technical details which make our implementation computationally efficient.
Figure 3: Fitting shape parameters with PyTorchGeoNodes. At every iteration, our genetic algorithm generates indiviudals of shape parameters and object poses. We parse 3D shapes from individual parameters using PyTorchGeoNodes and perform shape parameter optimization based on loss that evaluates how well the shape fits the input scene. In the next iteration, the optimized set of individuals is used to update the pool of individuals based on selection of the fittest principle.
@@ 4 Search Algorithm
We leverage our PyTorchGeoNodes framework to retrieve shape parameters and geometry of a target object in an RGB-D scan. We look for shape parameters ${\cal P}^{*}$ by minimizing an objective function that encourages the shape to fit the point cloud reconstructed from the images while lying on the floor:
${\cal L}({\cal P})=$ $\lambda_{\text{CD}}{\cal L}_{\text{CD}}({\text{PC}},{\text{Sh}}({\cal P}))+\lambda_{\text{G}}{\cal L}_{\text{G}}({\text{G}},{\text{Sh}}({\cal P}))+$
$\lambda_{\text{Fl}}{\cal L}_{\text{Fl}}({\text{PC}},{\text{Sh}}({\cal P})))\>.$ (1)
${\cal L}_{\text{CD}}({\text{PC}},{\text{Sh}}({\cal P}))$ measures the single-direction chamfer distance between PC, the point cloud reconstructed from the depth maps and object masks, and ${\text{Sh}}({\cal P})$, the shape generated by the shape program. We follow pre-processing from [1] which relies on sparse point cloud instance segmentation of the object and use a segmentation model [18] on images to obtain the object masks. We then back-project these masks to obtain a dense target point cloud. ${\cal L}_{\text{G}}({\text{G}},{\text{Sh}}({\cal P}))$ penalizes occlusions in the scene caused by ${\text{Sh}}({\cal P})$. G is an occlusion grid obtained by densely sampling 3D points along rays between the camera and surfaces for different views:
${\cal L}_{\text{G}}({\text{G}},{\text{Sh}}({\cal P}))=\sum_{p\in{\text{Sh}}({\cal P})}\begin{cases}\stackrel{{\scriptstyle\rightarrow}}{{{\text{CD}}}}(p,{\text{PC}})&\text{if }{\text{G}}(p)\\ 0&\text{otherwise},\end{cases}$ (2)
where $\stackrel{{\scriptstyle\rightarrow}}{{{\text{CD}}}}(p,{\text{PC}})$ is the distance of point $p$ to PC . ${\text{G}}(p)=\stackrel{{\scriptstyle\rightarrow}}{{{\text{CD}}}}(p,{\text{PC}})<\omega$ determines whether $p$ is on grid G. We set threshold $\omega=5cm$ in our experiments. Since G is generated using input views for the scene, this loss does not penalize points that are occluded in the scene, similar to render-and-compare objectives. The term ${\cal L}_{\text{Fl}}$ simply enforces the objects to stay on the floor plane:
${\cal L}_{\text{Fl}}({\text{PC}},{\text{Sh}}({\cal P})))=|{\text{Fl}}({\text{PC}})-{\text{Fl}}({\text{Sh}}({\cal P}))|_{2}\>,$ (3)